# Efficient Testing Implies Structured Symmetry

### v1

@@ 1.3 Technical Overview
(a) Tester $T$
(b) Simulator $\widetilde{T}_{j}$
(c) Simulator $\widetilde{T}_{j+1}$
Figure 1: Illustration of the supersimulator construction of Section 3. Plot (1(a)) depicts a deterministic 33-sample tester $T$ by the set of triples that cause it to output Accept. A one-way restriction of $T$ is shown in blue. Plot (1(b)) depicts a simulator in the sequence, with translucent regions indicating fractional estimates for $T$’s Accept region. Plot (1(c)) depicts the next simulator in the sequence, after an update has been performed based on the chosen restriction of $T$.
Our approach to proving Theorem 1.2, which states that efficient testing implies structured symmetry, will be to find a property $\mathcal{Q}$ such that $\mathcal{P}\subseteq\mathcal{Q}\subseteq\mathcal{P}_{\varepsilon}$ and membership in $\mathcal{Q}$ can be decided by a simple, structured approximation to a tester for $\mathcal{P}$. This approximation to the tester will itself be derived from simple, structured approximations to functions in $\mathcal{P}$ (or far from $\mathcal{P}$). For the latter (“simulating the oracle”) we use the complexity-theoretic regularity lemma of [TTV09], and for the former (“simulating the tester”) we use a recent enhancement of the lemma based on supersimulators [DT25].
At a high level, the complexity-theoretic regularity lemma states that any arbitrarily complex function $g:\mathcal{X}\to\{0,1\}$ has a low-complexity simulator $h:\mathcal{X}\to[0,1]$. By “simulator,” we mean that the error of $h$, namely ghg-h, is not too correlated with any function in a prespecified family $\mathcal{F}$ of distinguisher functions. Phrased differently, $g$ and $h$ are indistinguishable by $\mathcal{F}$. By “low-complexity,” we mean that $h$ is a weighted sum of a handful of functions from $\mathcal{F}$.
For example, if $\mathcal{X}=\{0,1\}^{n}$ and $\mathcal{F}$ comprises all Boolean circuits of size at most $s$, then the lemma states that any Boolean function $g:\{0,1\}^{n}\to\{0,1\}$ is indistinguishable from a circuit $h:\{0,1\}^{n}\to[0,1]$ of size [^3] We say a circuit $c$ with $n$ input bits, $m$ output bits, and $s$ logic gates computes a real-valued function $h:\{0,1\}^{n}\to[0,1]$ in binary if $h(x)=\sum_{i=1}^{m}c_{i}(x)/2^{i-1}$ for all inputs $x\in\{0,1\}^{n}$, where $c_{i}$ denotes the $i$th output bit of $c$. $O(s)$ with respect to all distinguisher circuits of size $s$. Notably, the simulator may be a constant factor larger than the distinguishers it is asked to fool. If we wish to construct simulators capable of fooling distinguishers that are larger than themselves, then we must instead use the stronger notion of supersimulators.
In what follows, we discuss the key steps of the proof in slightly more detail.
@@
First, we argue that we need not test arbitrarily complex functions $f:\{0,1\}^{n}\to\{0,1\}$ if the property of interest has a small tester $T$. Indeed, using the complexity-theoretic regularity lemma and a simple hybrid argument, we show that it suffices to test functions $\tilde{f}:\{0,1\}^{n}\to[0,1]$ that are computable by circuits slightly larger than $T$. Now, the oracle simulator returns modeled labels $\tilde{y}_{i}$ sampled from the Bernoulli distribution $\mathcal{B}(\tilde{f}(x_{i}))$, rather than real labels $y_{i}=f(x_{i})$. To the tester, however, they look no different.
@@
Next, consider two $m$-sample testers $T$ and $\widetilde{T}$. Suppose that they both receive labels generated from a circuit $\tilde{f}$ that is slightly larger than both testers, and that this circuit happens to be $\{0,1\}$-valued. (Technically, $\tilde{f}$ may take fractional values, but we address this challenge in the full version of the proof.) We show that $T$ and $\widetilde{T}$ have similar probabilities of outputting Accept if $T$ and $\widetilde{T}$ are indistinguishable by a certain function derived from $\tilde{f}$, called $\tilde{f}^{\prime}$. This function receives $m$ labeled samples and checks whether they are all consistent with $\tilde{f}$:
$\tilde{f}^{\prime}(x,y)=\bm{1}\bigl[\forall i\in[m],\,y_{i}=\tilde{f}(x_{i})\bigr].$
Ideally, we would like to apply the complexity-theoretic regularity lemma once more to construct a small simulator $\widetilde{T}$ that fools $\tilde{f}^{\prime}$ for any circuit $\tilde{f}$ that is slightly larger than $T$ and $\widetilde{T}$. Indeed, if we could do this, then we would be done. The lemma would decompose $\widetilde{T}$ into a simple weighted sum of functions $\tilde{f}_{1}^{\prime},\ldots,f_{k}^{\prime}$, and taking all intersections of level sets of the low-complexity functions $\tilde{f}_{1},\ldots,\tilde{f}_{k}$ would give us the low-complexity, symmetric partition that we desire.
@@
The discerning reader will notice the flaw in the preceding argument: we require the simulator $\widetilde{T}$ to fool distinguishers $\tilde{f}^{\prime}$ that are slightly more complex than itself, which the complexity-theoretic regularity lemma cannot provide! To overcome this obstacle, we apply supersimulators, which are a powerful strengthening of the lemma that constructs simulators capable of fooling significantly more complex distinguishers.
At a high level, the supersimulator construction we employ is iterative, generating a sequence of approximations $\widetilde{T}_{1},\widetilde{T}_{2},\ldots$ to the original tester $T$. At the $j$th step, we derive $\widetilde{T}_{j+1}$ from $\widetilde{T}_{j}$ by first choosing a function $\tilde{f}_{j}$ slightly larger than $\widetilde{T}_{j}$ and then adding an appropriate multiple of $\tilde{f}_{j}^{\prime}$ to $\widetilde{T}_{j}$. Specifically, the function $\tilde{f}_{j}$ will a combination of a handful of one-way restrictions of $\widetilde{T}_{j}$, which are obtained by hard-wiring all but one input to the current simulator $\widetilde{T}_{j}$. We illustrate the construction in Figure 1. We discuss supersimulators further in Section 2, give a full proof of Theorems 1.2 and 1.3 in Section 3, and prove our various extensions in Section 4.
@@ 1.4 Related Work
There is a large body of work devoted to decomposing complex objects into their structured and pseudorandom components. An early result in this space is Szemerédi’s regularity lemma [Sze75]. The result roughly states that any large, dense graph can be split into a (rather large) handful of parts between which the graph’s edges are distributed pseudorandomly. There is a vast literature on the regularity lemma and its variants, producing breakthrough results in pseudorandomness and additive combinatorics to this day [KM23, JLL+25]. For more, see the surveys [Tao07, Zha23].
An especially relevant result in this field is the complexity-theoretic regularity lemma of [TTV09], which gave a unified perspective on such disparate topics as the Frieze-Kannan weak regularity lemma for graphs [FK96, FK99], Impagliazzo’s hardcore lemma [Imp95], and the dense model theorem [GT08, TZ08, RTTV08]. It also led to a deeper understanding of computational entropy [VZ12, VZ13, Zhe14] and techniques for leakage simulation in cryptography [JP14].
@@
The connection between these modern tools and older notions of regularity was made explicit by [DLLT23, CDV24], and this point of view has proven to be fruitful. Indeed, these works and their sequelae have led to new insights into graph regularity, hardcore set construction, dense models, omniprediction, computational hardness and entropy, and the computational indistinguishability of product distributions [MPV25, CGKR25, HV25, DT25]. The present work similarly builds on the ideas from this line of work, but now with a view toward property testing. In particular, supersimulation [DT25], which is our main technical tool, is closely related to both outcome indistinguishability and multicalibration.
@@
In the present paper, we do not work with dense graphs, but instead test properties of a certain class of dense distributions defined by Boolean functions. As already discussed, in this context, the closest related work to ours is the characterization of constant-sample testability in terms of constant-part symmetry, due to [BY19]. Even earlier, [KS08, Sud10] pioneered the idea that the symmetries of a property play a central role in understanding its testability.
Our work concerns the computational complexity of property testing, which has recently received renewed interest [FPR26]. For the related but simpler task of efficiently determining which of two distributions $\mathcal{D}_{0}$ vs $\mathcal{D}_{1}$ generated an observed set of $m$ samples, the recent work of [MPV25] used multicalibration to argue that densities on a low-complexity partition of the domain contain all the relevant information (see also [DT25]). Our results have a similar flavor, but we test for membership in sets $\mathcal{P}$ vs $\neg\mathcal{P}_{\varepsilon}$ that may contain doubly exponentially many distributions, rather than just one each.

### v2

@@ 1.3 Technical Overview
The results of this paper largely rely on a technical tool of independent interest, which we call supersimulators. Although the existence of supersimulators is implicit in prior work [DKR+21], the present work is the first to explore their complexity-theoretic implications. In Section 1.3.1, we motivate supersimulators, discuss their connection to graph regularity, and explain how, surprisingly, they circumvent known lower bounds to improving the complexity-theoretic regularity lemma of [TTV09]. In Section 1.3.2, we explain how they are used in the proof of our first main result, Theorem 1.2.
1.3.1 Supersimulators and the Regularity Lemma
As previously discussed, the Szemerédi regularity lemma [Sze75] is a cornerstone result in graph theory that splits any large, dense graph—no matter how complex—into a small number of parts between which the graph’s edges are distributed pseudorandomly. Although Szemerédi’s regularity lemma spawned many variants, these versions were still fairly specialized (e.g. to cut sizes in graphs, or Fourier uniformity in vector spaces over finite fields) until the arrival of the complexity-theoretic regularity lemma [TTV09], which generalized and abstracted the concept of regularity by considering indistinguishability with respect to an arbitrary collection of distinguisher functions defined on an arbitrary domain.
For example, consider a setting especially relevant to applications in complexity theory and cryptography: distinguishers computable by size-bounded Boolean circuits that receive a uniformly random input from $\{0,1\}^{n}$. In this setting, the regularity lemma states that every randomized Boolean function of arbitrary complexity can be simulated by a randomized circuit of size $O(s)$ that fools all distinguishers of size at most $s$, for any $s$.
Lemma 1.19 (Special Case of Theorem 1.1 of [TTV09]).
For all target functions $g:\{0,1\}^{n}\to[0,1]$, sizes $s\in\mathbb{N}$, and error tolerances $\varepsilon>0$, there exists a simulator $h:\{0,1\}^{n}\to[0,1]$ of circuit size [^7] We say a circuit $c$ with $n$ input bits, $m$ output bits, and $s$ logic gates computes a real-valued function $h:\{0,1\}^{n}\to[0,1]$ in binary if $h(x)=\sum_{i=1}^{m}c_{i}(x)/2^{i-1}$ for all inputs $x\in\{0,1\}^{n}$, where $c_{i}$ denotes the $i$th output bit of $c$. at most $O(s/\varepsilon^{2})$ such that for all distinguishers $A:\{0,1\}^{n+1}\to\{0,1\}$ of size at most $s$,
[…]
Lemma 1.19 ensures the existence of a simulator that is at most a constant factor larger than the distinguishers, but makes no guarantees about the existence of smaller simulators. At least three different arguments in the literature, two in the original paper [TTV09] and one in a subsequent work [CCL18], assert that this complexity gap is inevitable, and the simulator must always be allowed to be larger than the distinguishers that it is asked to fool. For more on these barriers, see the discussion of related work in Section 1.4.
Surprisingly, all three lower bound arguments can be circumvented by relaxing the requirement that the result hold for all size bounds $s$. By instead allowing the choice of $s$ to depend on the target function, while remaining below an upper bound independent of the target function, it is possible to design simulators that fool families of distinguishers far more powerful than themselves. In this work, we will refer to such an object as a supersimulator. As our main results demonstrate, this relaxed quantification in which $s$ depends on $g$ is still useful for applications.
The existence of supersimulators is implicit in the code-access outcome indistinguishability construction of [DKR+21], in which the distinguishers to be fooled have access to the code of the simulator (and are therefore perforce larger). Thus, the following statement can be viewed as a rephrasing of Theorem 5.10 of [DKR+21], unencumbered by questions of the distinguishers’ access to the code of the simulator. A related result appeared even earlier, albeit in a slightly different computational model, in the study of key-agreement in cryptography [HNO+18].
Lemma 1.20 (Supersimulators, Special Case [DKR+21]).
For all $g:\{0,1\}^{n}\to\{0,1\}$, $k\in\mathbb{N}$, and $\varepsilon>0$, there exists a size bound $s\in[n,n^{k^{O(1/\varepsilon^{2})}}]$ and a simulator $h:\{0,1\}^{n}\to[0,1]$ of size at most $s$ such that for all distinguishers $A:\{0,1\}^{n+1}\to\{0,1\}$ of size at most $s^{k}$,
[…]
To understand the statement of Lemma 1.20, suppose that $k=100$, $\varepsilon=1/10$, and the average-case complexity of $g$ is superpolynomial in $n$. In this case, one cannot hope to approximate $g$ accurately with a function $h$ of complexity $s=\mathrm{poly}(n)$, but Lemma 1.20 nevertheless guarantees that there exists a simulator of this size fooling all distinguishers of size up to $s^{100}$, well beyond the capabilities of the simulator provided by Lemma 1.19. The particular choice of $s$ may depend on $g$, but it never leaves the interval $n,n^{10^{300}}]$, say, which is fixed and independent of $g$. For clarity, we state this concrete version separately, as a corollary:
Corollary 1.21.
For all $g:\{0,1\}^{n}\to\{0,1\}$, $k\in\mathbb{N}$, and $\varepsilon>0$, there exists $s=\mathrm{poly}(n)$ and $h:\{0,1\}^{n}\to[0,1]$ of size at most $s$ such that for all $A:\{0,1\}^{n+1}\to\{0,1\}$ of size at most $s^{100}$,
[…]
While we have stated both Lemmas 1.19 and 1.20 in the setting of size-bounded Boolean circuits on the domain $\mathcal{X}=\{0,1\}^{n}$, both results generalize to arbitrary distinguishers families on an arbitrary domain. We call this the abstract setting and review it further in Section 2.
In Appendix A, we present two proofs of Lemma 1.20. The first more closely resembles the proof in [DKR+21], as well as the proof of the original complexity-theoretic regularity lemma in [TTV09], but we use a slightly modified projection operation which will be important for our downstream applications.
Our second proof is based on an iteration technique from the graph regularity literature, in which strong notions of regularity are achieved by iterating cheaper constructions with a shrinking sequence of error parameters. In graph theory, this technique has been used to achieve state-of-the-art quantitative bounds for applications to graph removal [Fox11, CF12] and to establish relationships between existing notions of regularity [RS10]. In our setting, we use it to increase the power of the simulators, allowing them to fool distinguishers whose complexity surpasses their own by any specified “growth function.” This proof technique also has the advantage of a slightly stronger theorem statement, which incorporates a shrinking error tolerance that decays as a function of the simulator size. We will not, however, require this additional power for our applications to property testing.
1.3.2 Finding Structured Symmetry
In this section, we sketch the proof of our first main result, Theorem 1.2. This result states that any property $\mathcal{P}$ with an efficient tester $T$ is roughly equivalent to a property $\mathcal{Q}$ with structured symmetry. Here, “roughly equivalent” means that $\mathcal{P}\subseteq\mathcal{Q}\subseteq\mathcal{P}_{\varepsilon}$.
Our approach will be to construct a simulator $\widetilde{T}$ for $T$ that has a particularly simple form, which we will then use to define the property $\mathcal{Q}$. Specifically, we shall say that a function $f$ has the property $\mathcal{Q}$ if the tester $\widetilde{T}$ is more likely to output Accept than Reject when it receives $m$ independent inputs $(x_{i},y_{i})$ labeled by $f$. Here, $x_{i}\sim\{0,1\}^{n}$ is a uniform $n$-bit string and $y_{i}=f(x_{i})$. We will show that $\mathcal{P}\subseteq\mathcal{Q}\subseteq\mathcal{P}_{\varepsilon}$ by using the fact that $\widetilde{T}$ is a simulator for a tester for $\mathcal{P}$. Similarly, we will argue that $\mathcal{Q}$ has structured symmetry by reasoning about the simple form of $\widetilde{T}$.
[…]
Step 0: Notation
Before proceeding in more detail, it will be helpful to review the relevant lemmas using slightly simpler notation than before. Indeed, the complexity-theoretic regularity lemma can be viewed as stating that every function $g:\mathcal{X}\to\{0,1\}$ has a low-complexity simulator $h:\mathcal{X}\to[0,1]$. By “simulator,” we mean that the error of $h$, namely ghg-h, is not too correlated with any function in a prespecified family $\mathcal{F}$ of distinguisher functions:
[…]
By “low-complexity,” we mean that $h$ is a weighted sum of a handful of functions from $\mathcal{F}$:
$h(x)=h_{k}(x)=\Bigl[\delta\cdot\bigl(f_{1}(x)+\cdots+f_{k}(x)\bigr)\Bigr]_{0}^{1}$
for some $f_{1},\ldots,f_{k}\in\pm\mathcal{F}$, where $\cdot]_{0}^{1}$ denotes projection onto the interval $[0,1]$. The supersimulators lemma replaces the fixed family $\mathcal{F}$ with a variable family $\mathcal{F}(h)$ which grows steadily more complex as $h$ grows in circuit size, meaning that for each index $j\in[k]$, we have $f_{j}\in\pm\mathcal{F}(h_{j-1})$, where $h_{j-1}$ is defined analogously to $h_{k}$ above.
In what follows, we model the tester as a function $T:(\mathcal{X}\times\{0,1\})^{m}\to\{0,1\}$ which takes as input $m$ labeled samples $(x_{i},y_{i})\in\mathcal{X}\times\{0,1\}$ and outputs either 0 or $1$, corresponding to Reject or Accept. For simplicity, suppose for now that $T$ is deterministic. (We will handle randomized testers in the full version of the proof.) Also note that $\mathcal{X}=\{0,1\}^{n}$.
@@ 1.3.2 Finding Structured Symmetry
First, we argue that if the property $\mathcal{P}$ of interest has a tester $T$ computable by a small circuit of size $s$, then we need not consider arbitrarily complex labeling functions $f:\{0,1\}^{n}\to\{0,1\}$. Using the complexity-theoretic regularity lemma and a simple hybrid argument, we show that it suffices to study functions $\tilde{f}:\{0,1\}^{n}\to[0,1]$ that are computable by circuits of size $O(s)$, which is to say at most a constant factor larger than $T$. Now, the oracle simulator returns modeled labels $\tilde{y}_{i}$ sampled from the Bernoulli distribution $\mathcal{B}(\tilde{f}(x_{i}))$, rather than real labels $y_{i}=f(x_{i})$, but they look no different to the tester. Slightly more formally, we suppose that $\tilde{f}$ is a $(\mathcal{F},\delta)$-regular simulator for $f$ with respect to the class $\mathcal{F}$ of one-way restrictions of $T$, which are functions obtained by hardwiring all but one input to $T$ (note that if $T$ has a circuit of size at most $s$, then so do functions in $\mathcal{F}$). Under this assumption on $\tilde{f}$, we show that
$\Bigl\lvert\operatorname*{\mathbb{E}}\bigl[T(x,y)\bigr]-\operatorname*{\mathbb{E}}\bigl[T(x,\tilde{y})\bigr]\Bigr\rvert\leq 2m\delta.$
By the complexity-theoretic regularity lemma, there always exists such a simulator $\tilde{f}$ of circuit size $O(s)$, which, notably, is at most slightly larger than the circuit complexity of $T$.
@@
Next, consider two $m$-sample testers $T$ and $\widetilde{T}$. Suppose that they both receive labels generated from a circuit $\tilde{f}$ that is slightly larger than both testers, as suggested by the construction at the end of the previous step. Suppose also that $\tilde{f}$ happens to be $\{0,1\}$-valued. (We will address the challenge of fractional values in the full version of the proof.) We show that $T$ and $\widetilde{T}$ have similar probabilities of outputting Accept if $T$ and $\widetilde{T}$ are indistinguishable by a certain function derived from $\tilde{f}$, called $\tilde{f}^{\prime}$. This function receives $m$ labeled samples and checks whether they are all consistent with $\tilde{f}$:
$\tilde{f}^{\prime}(x,y)=\bm{1}\bigl[\forall i\in[m],\,y_{i}=\tilde{f}(x_{i})\bigr].$ (1.22)
Slightly more formally, we show that if $\widetilde{T}$ is a simulator for $T$ with error $\gamma$ with respect to the class of functions $\tilde{f}^{\prime}$, where $\tilde{f}$ is any circuit of size at most $)O(\max\{\mathrm{size}(T),\mathrm{size}(\widetilde{T}$)\}), then
$\Bigl\lvert\operatorname*{\mathbb{E}}\bigl[T(x,\tilde{y})\bigr]-\operatorname*{\mathbb{E}}\bigl[\widetilde{T}(x,\tilde{y})\bigr]\Bigr\rvert\leq 2^{m}\gamma.$
(Recall from the previous step that $\tilde{y}_{i}=\tilde{f}(x_{i})$ are modeled labels generated by $\tilde{f}$.)
Ideally, we would like to apply the complexity-theoretic regularity lemma once more to construct a small simulator $\widetilde{T}$ that fools $\tilde{f}^{\prime}$ for any circuit $\tilde{f}$ that is slightly larger than $T$ and $\widetilde{T}$. This includes, in particular, the function $\tilde{f}$ constructed in the previous step, applied to both $T$ and $\widetilde{T}$. Indeed, if we could do this, then we would be done. We would simply define $\mathcal{Q}$ to be the property of a function $f$ such that $f\in\mathcal{Q}$ iff the tester $\widetilde{T}$ is more likely to accept than reject on labels generated by $f$. We would be able to argue that $\mathcal{P}\subseteq\mathcal{Q}\subseteq\mathcal{P}_{\varepsilon}$ using the relationship between the output of $T$ and $\widetilde{T}$ on labels $y$ generated by $f$:
$\operatorname*{\mathbb{E}}\bigl[T(x,y)\bigr]\approx_{2m\delta}\operatorname*{\mathbb{E}}\bigl[T(x,\tilde{y})\bigr]\approx_{2^{m}\gamma}\operatorname*{\mathbb{E}}\bigl[\widetilde{T}(x,\tilde{y})\bigr]\approx_{2m\delta}\operatorname*{\mathbb{E}}\bigl[\widetilde{T}(x,y)\bigr].$
Moreover, setting $\delta=O(1/m)$ and $\gamma=O(1/2^{m})$, the lemma would decompose $\widetilde{T}$ into a simple weighted sum of $\mathrm{poly}(1/\gamma)=2^{O(m)}$ functions $\tilde{f}_{1}^{\prime},\ldots,\tilde{f}_{k}^{\prime}$. Taking all intersections of level sets of the low-complexity functions $\tilde{f}_{1},\ldots,\tilde{f}_{k}$ would give us the efficiently computable, symmetric partition into $2^{2^{O(m)}}$ parts that we desire.
@@
There is a flaw in the preceding argument, as stated: We require the simulator $\widetilde{T}$ to fool distinguishers $\tilde{f}^{\prime}$ that are slightly more complex than itself, which the complexity-theoretic regularity lemma cannot provide! In the full proof, we overcome this obstacle using supersimulators.
Unpacking definitions, we see that the supersimulator construction we employ is iterative, generating a sequence of approximations $\widetilde{T}_{1},\widetilde{T}_{2},\ldots$ to the original tester $T$. At the $j$th step, we derive $\widetilde{T}_{j+1}$ from $\widetilde{T}_{j}$ by first choosing a function $\tilde{f}_{j}$ slightly larger than $\widetilde{T}_{j}$ and then adding an appropriate multiple of $\tilde{f}_{j}^{\prime}$ to $\widetilde{T}_{j}$. Specifically, the function $\tilde{f}_{j}:\mathcal{X}\to[0,1]$ will be a combination of several one-way restrictions of $\widetilde{T}_{j}$, which are obtained by hard-wiring all but one input to the current simulator $\widetilde{T}_{j}$, and the function $\tilde{f}_{j}^{\prime}:(\mathcal{X}\times\{0,1\})^{m}\to\{0,1\}$ is defined with respect to $\tilde{f}_{j}$ as per equation (1.22).
@@ 1.4 Related Work
There is a large body of work devoted to decomposing complex objects into their structured and pseudorandom components. As already discussed, an early result in this space is Szemerédi’s regularity lemma [Sze75]. There is a vast literature on the regularity lemma and its variants, producing breakthrough results in pseudorandomness and additive combinatorics to this day [KM23, JLL+25]. For more, see the surveys [Tao07, Zha23].
The complexity-theoretic regularity lemma of [TTV09] gave a unified perspective on such disparate topics as the Frieze-Kannan weak regularity lemma for graphs [FK96, FK99], Impagliazzo’s hardcore lemma [Imp95], and the dense model theorem [GT08, TZ08, RTTV08]. It also led to a deeper understanding of computational entropy [VZ12, VZ13, Zhe14] and techniques for leakage simulation and key-agreement in cryptography [JP14, CCL18, HNO+18]. Similar ideas to those of [TTV09] also appeared in [Fel12, DDFS12].
The Regularity Barrier
In the Boolean circuit case, the regularity lemma (Lemma 1.19) asserts that for any size bound $s\in\mathbb{N}$, there exists a simulator $h$ of size $s^{\prime}=O(s)$ that fools distinguishers of size $s<s^{\prime}$. The question of the existence of size-$s^{\prime}$ simulators $h$ that fool distinguishers of size $s\gg s^{\prime}$ has already been considered by multiple works [TTV09, CCL18], with negative results.
First, Remark 1.6 of [TTV09] constructed two counterexamples ruling out the possibility of any version of Lemma 1.19 with $s\geq(ns^{\prime})^{1+\Omega(1)}$. One of their examples involves a target function $g:\{0,1\}\to\{0,1\}$ of complexity $\tilde{O}(ns^{\prime})$ sampled from a family of $O(s^{\prime}\log s^{\prime})$-wise independent hash functions. Using a Chernoff-like concentration inequality for $k$-wise independence (e.g. Problem 3.8 of [Vad12]), they show that with high probability over the choice of $g$, every function $h$ of complexity at most $s^{\prime}$ has correlation at most $1/10$ with $g$. Consequently, one cannot hope for $h$ to fool the distinguisher $f=g$ of size $s=\tilde{O}(ns^{\prime})$.
Crucially, in the aforementioned counterexample, the target function $g$ depends on the simulator size $s^{\prime}$ (for example, in one standard construction of the hash family, $g$ would be a polynomial whose degree grows with $s^{\prime}$). The same is true of their other counterexample, which is based on a black-box application of randomness extractors for efficiently sampleable distributions with high min-entropy. Therefore, these counterexamples do not rule out supersimulators (Lemma 1.20), which only guarantee that for all $g$, there exists at least one “good” simulator size $s^{\prime}$ in a bounded interval.
The subsequent work of [CCL18] provided an even stronger lower bound on the simulator size, under additional assumptions on its structure. Specifically, they argue that any “black-box” $(\mathcal{F},\varepsilon)$-regular simulator $h$ for $g$ under $\mathcal{D}$ must make $\Omega(1/\varepsilon^{2})$ oracle calls to functions in $\mathcal{F}$. In the Boolean circuit case, this result may lead one to believe that simulators of size $s^{\prime}$ can only hope to fool substantially smaller distinguishers, and not even those of slightly smaller sizes $s\in[\varepsilon^{2}s^{\prime},s^{\prime}]$. From one point of view, Lemma 1.20 shows that in the context of Boolean circuits, removing this structural requirement on $h$ impacts the range of attainable simulator and distinguisher sizes.
@@
The connection between these modern machine learning tools and older notions of regularity was made explicit by [DLLT23, CDV24], and this point of view has proven to be fruitful. Indeed, these works and their sequelae have led to new insights into graph regularity, hardcore set construction, dense models, omniprediction, computational hardness and entropy, and the computational indistinguishability of product distributions [MPV25, CGKR25, HV25]. The present work similarly builds on the ideas from this line of work, but now with a view toward property testing. In particular, supersimulation, which is our main technical tool, is closely related to both outcome indistinguishability and multicalibration.
Calibrated Multiaccuracy
The idea behind our proof of Theorem 1.17, which concerns the computational indistinguishability of product distributions as in [MPV25], is to replace the use of multicalibration (a sometimes costly strengthening of the regularity lemma) with calibrated multiaccuracy (a lightweight strengthening of the regularity lemma). The use of this idea to achieve quantitative gains in downstream applications has been a theme of multiple recent and concurrent works.
One such work shows that the concept of omniprediction in machine learning can be achieved via calibrated multiaccuracy [GHK+23] (see also [DLLT23]), which had previously been shown to be achievable via multicalibration [GKR+22]. Another work shows that for hardcore set construction, optimal density can be achieved via calibrated multiaccuracy [CGKR25], which had previously been shown to be achievable via multicalibration [CDV24], and even earlier with suboptimal density via the original regularity lemma [TTV09]. Most recently, [HV25] showed that calibrated multiaccuracy can be used to extend and improve the efficiency of prior regularity-based characterizations of computational notions of entropy [VZ12, VZ13, Zhe14, CDV24].
@@
In the present paper, we focus not only on properties of dense graphs, but also on properties of a certain class of dense distributions defined by Boolean functions. As already discussed, in this context, the closest related work to ours is the characterization of constant-sample testability in terms of constant-part symmetry, due to [BY19]. Even earlier, [KS08, Sud10] pioneered the idea that the symmetries of a property play a central role in understanding its testability. Of course, one difference is that our work concerns the computational complexity of property testing, which has recently received renewed interest [FPR26].
Private Graph Algorithms
There have been several works designing differentially private algorithms for approximating the cut function of a graph $G=(V,E)$ [GRU12, AU19, EKKL20, LUZ24, ZLU25, ACD+25]. This function takes as input two vertex subsets $S,T\subseteq V$, and outputs their edge count $e_{G}(S,T)$. Our algorithm for privately compute regularity templates has similar (but, technically speaking, incomparable) goals and guarantees. While state-of-the-art algorithms for privately approximating the cut function generally take $\mathrm{poly}(n)$ or $\exp(n)$ time and achieve $O(n^{c})$ error for $c<2$, our algorithm for compute regularity templates allows one to approximate cut sizes (without location information) with $\varepsilon n^{2}$ by inspecting only a constant $O_{\varepsilon}(1)$ number of vertices of the graph. Phrased differently, our algorithm computes a coarser approximation of the graph, but does so much more efficiently—indeed, more efficiently than one might expect to be possible.
1.5 Paper Organization
In Section 2, we cover preliminaries on property testing, regularity, and supersimulators. In Section 3, we prove our main results on efficient property testing, Theorems 1.2 and 1.3, as well as our result on consistency counting, Theorem 1.5, and our extension to dense distribution testing, Theorem 1.6. In Section 4, we present our results on node-level private graph testing, Theorems 1.12 and 1.13, as well as the related function testing results, Theorems 1.15 and 1.16. In Section 5, we prove our results on the computational indistinguishability of product distributions: Theorems 1.17 and 1.18. In Appendix A, we present the constructions of supersimulators on which we rely. These imply the version stated in Lemma 1.20 as a special case.